How many solutions are there to the equation x1+x2+x3+x4+x5+x6= 25 where the xi(for i= 1,2,3,4,5,6) are non-negative integers?
1
Expert's answer
2020-11-24T08:09:38-0500

k-combination with repetitions, or multisubset of size "k"  from a set "S"  is given by a sequence of "k"  not necessarily distinct elements of "S", where order is not taken into account: two sequences define the same multiset if one can be obtained from the other by permuting the terms. Associate an index to each element of "S"  and think of the elements of "S"  as types of objects, then we can let "{\displaystyle x_{i}}" denote the number of elements of type "i" in a multisubset. The number of multisubsets of size "k"  is then the number of nonnegative integer solutions of the Diophantine equation:


"x_1+x_2+...+x_n=k"


If "S"  has "n"  elements, the number of such "k"-multisubsets is denoted by "({n \choose k})". This expression, "n"  multichoose "k", can also be given in terms of binomial coefficients:


"({n \choose k})={n+k-1\choose k}".


In our case, the number of nonnegative solutions of the equation


"x_1+x_2+x_3+x_4+x_5+x_6= 25"


is equal to

"({6 \choose 25})={30\choose 25}=\frac{30!}{25!\cdot5!}=\frac{30\cdot 29\cdot 28\cdot 27\cdot 26}{5\cdot 4\cdot 3\cdot 2}=142,506"